# VidHal: Benchmarking Temporal Hallucinations in Vision LLMs

### v1

@@ 5 Experiments
Model Release Date Visual Encoder LLM #Params #Frames Accuracy NDCG
Naive Relative
Baseline
Random - - - - - 0.326 0.505 0.480
Open-Sourced Models
VideoChat2 [28] Apr. 2024 UMT-L Vicuna 7B 16 0.410 0.490 0.573
Mistral 7B 16 0.524 0.348 0.579
Phi3 3.8B 16 0.468 0.552 0.522
mPLUG-Owl3 [55] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
LLaVA-NeXT-Video [62] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
Qwen1.5 32B 32 0.663 0.641 0.747
VideoLLaMA2 [9] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
Qwen2 72B 8 0.647 0.787 0.760
Proprietary Models
GPT-4o [41] May 2024 - - 1fps 0.772 0.840 0.826
Gemini-1.5 [44] Feb. 2024 Flash - 1fps 0.657 0.738 0.745
Pro - 1fps 0.671 0.765 0.753
Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task.
@@ 5.1 Experimental Settings
Models. We evaluated eleven VLLMs from six different model families, including four open-source models: VideoLLaMA2 [9], VideoChat2 [28], mPLUG-Owl3 [55] and LLaVA-NeXT-Video [62], and two proprietary models: GPT-4o [41] and Gemini-1.5 [44]. These models represent a wide variety of architectural designs and training paradigms. Additionally, we included a random baseline that selects and ranks candidate options randomly.
Implementation Details. All experiments were conducted using four NVIDIA A100 40GB GPUs. The input captions in $\mathcal{Y}^{i}$ are displayed to each model in a randomized order, with a predefined randomization seed applied consistently across all experiments. This operation aims to prevent the model from consistently making the same prediction regardless of the caption orders. We adhered to the inference and model hyperparameters outlined in the respective original models, and employed greedy decoding during generation for a fair comparison.
@@ 5.2 Overall Results
We present the overall results of representative VLLMs in Table 2 across both MCQA and caption ordering tasks. We make three key observations from this table: I) Proprietary models demonstrate superior results compared to open-sourced models. In particular, GPT-4o achieves the best performance on all tasks, surpassing other models by significant margins. II) Larger VLLMs generally outperform smaller ones in both tasks. This result is supported by the comparison of different LLM bases for the VideoLLaMA2 and LLaVA-NeXT-Video models. III) The caption ordering task poses greater difficulty for current VLLMs than MCQA, evidenced by the larger performance margins between the VLLM models and the random baseline. Notably, the VideoChat2 models show slight to no improvement over the random baseline across both caption ordering tasks. This indicates that current VLLMs greatly suffer from poor fine-grained video understanding and are inclined to generate hallucinations.
Figure 6: Distribution of results of VLLMs across varied input caption orders for the three evaluation tasks.
@@ 5.3 Ablation Studies
Input Order Sensitivity
To assess the robustness of VLLM responses to the order of displayed captions, we conducted additional experiments by evaluating three VLLMs using a fixed static display order across all instances. We repeated this process across all different permutations of input caption order, presenting the results of these models in Figure 6. We observe that the performance of these VLLMs is highly sensitive to the order in which captions are displayed, reflected by their varying results across different order permutations. This instability intensifies with smaller model sizes, with VideoLLaMA2 (7B) showing the highest variance in evaluation results and VideoLLaMA2 (72B) the lowest. Our findings suggest that VLLMs may be particularly vulnerable to input caption order, potentially confounding their performance.
Figure 7: (Top) Averaged consensus score of each respective model, and (Bottom) consensus between each model pairs. Average represents the average consensus score across all VLLMs, with Gemini-1.5 Pro achieving the highest.
Model Correct Incorrect Overall
Naive Caption Ordering
VideoLLaMA2 (7B) 0.889 0.940 0.914
LLaVA-NeXT-Video (32B) 0.727 0.709 0.718
VideoLLaMA2 (72B) 0.726 0.568 0.642
Relative Caption Ordering
VideoLLaMA2 (7B) 0.799 0.781 0.790
LLaVA-NeXT-Video (32B) 0.710 0.579 0.641
VideoLLaMA2 (72B) 0.670 0.444 0.545
Table 3: Overlapping ratios of model predictions under single-frame inputs and full-videos for correctly predicted, incorrectly predicted, and overall of both cases.
Model Consensus
We further assess the reliability of VLLM responses by measuring their consensus on the MCQA task, defined as the proportion of shared correct predictions across correctly answered instances between model pairs. As shown in Figure 7, advanced VLLMs typically show higher consensus with other models sharing similar performance. In contrast, smaller models display less consistency in reasoning, with VideoChat2 demonstrating especially low response agreement among the smaller VLLMs.
@@
Image Prior Reliance
Prior research reveals that VLLMs often leverage image priors for reasoning [23, 3], while ignoring the important spatiotemporal features. This phenomenon is exemplified by a few frames having a dominant influence on response generation. To investigate whether the image prior issue influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [46] to extract the most salient frame $v^{i}$ from $V^{i}$. Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input. The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings. As illustrated in Table 3, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions. This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
5.4 In-Depth Analysis
We further leverage three quantitative metrics to identify other deficiencies in VLLM responses.
@@ 5.4 In-Depth Analysis
  • Hallucination Misalignment (HM) exposes the tendency in visual-language models to favor higher-hallucination captions over lower-hallucination ones in relative caption ranking tasks. We introduce two sub-metrics to detect such misalignments: $HM_{3\rightarrow 1}$ and $HM_{3\rightarrow 2}$, representing the proportion of cases where the VLLM selects the most hallucinatory caption over the anchor and the lower-hallucination caption, respectively. The general form for both terms can be expressed formally as follows:
@@
Figures 8 presents the IRR and RR scores for all evaluated models. Despite being able to formulate responses with the correct structure, many VLLMs tend to produce the same caption order regardless of $V^{i}$, as reflected by their high RR scores. This is particularly evident in models with limited capacity, such as those equipped with LLMs in the 7B parameter range.
Advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9. However, all evaluated models possess greater $HM_{3\rightarrow 2}$ scores, indicating difficulty in differentiating between two hallucinatory captions with varying levels of hallucination. This gap is especially pronounced in advanced VLLMs like GPT-4o, with $HM_{3\rightarrow 2}$ being significantly closer to the random baseline compared to $HM_{3\rightarrow 1}$. These results reinforce that gaps in nuanced video reasoning may potentially drive hallucinatory behavior in VLLMs, a challenge overlooked by existing paired ⟨positive, hallucinatory⟩-based evaluations [29, 50].
@@ 6 Conclusion
In this work, we introduce the VidHal benchmark to address gaps in the video-based hallucination evaluation of VLLMs. VidHal features video instances spanning five temporal aspects. Additionally, we propose a novel caption ordering evaluation task to probe the fine-grained video understanding capabilities of VLLMs. We conduct extensive experiments on VidHal through the evaluation of eleven VLLMs, exposing their limitations in unexpected hallucination generation. Our empirical results shed light on several promising directions for future work: e.g., fine-tuning models with human feedback to reduce undesired regurgitation responses, and countering single-frame priors for improving temporal reasoning. These advancements, on the other hand, will mitigate the hallucination problem in video-based VLLMs, thereby improving their robustness for real-world video understanding applications.
A Benchmark Construction Details
A.1 Dataset Statistics
Figure 11 illustrates the distribution of public dataset sources contributing to the visual instances in VidHal. Additionally, Figures 12 and 13 depict the distribution of temporal aspects across VidHal and the ground truth answers for the MCQA and caption ordering tasks, respectively. One can observe that both temporal aspects and ground truth options are uniformly distributed across our benchmark.
Figure 11: Distribution of visual instances in VidHal by public dataset source, categorized across the five temporal aspects.
Figure 12: Distribution of temporal aspects and subaspects in VidHal.
Figure 13: Distribution of (Left) correct answer options for the MCQA task and (Right) optimal option orders for the caption ordering task.
A.2 Dataset Development Pipeline
Figure 14: Specific skills and corresponding questions from the Perception Test dataset chosen for VidHal instance selection, with the matched aspects indicated in brackets.
Figure 15: Evaluation tasks in MVBench aligned with temporal aspects in VidHal, categorized by aspect.
Visual Instance Selection
To ensure a rich coverage of temporal aspects and visual diversity, we methodically selected video instances from four public datasets: TempCompass [38], Perception Test [43], MVBench [28], and AutoEval Video [6]. Given the unique characteristics of each dataset, we outline the specific guidelines adopted for each dataset below:
  • TempCompass encompasses five temporal aspects: Action, Speed, Direction, Event Order, and Attribute Change. As most of these aspects align with those chosen to construct VidHal, we retain all video instances except those related to speed. TempCompass includes four evaluation tasks: MCQA, Yes/No QA, caption matching, and caption generation. Given the conciseness of captions in the latter two tasks, their information can often be subsumed within the more detailed QA-based annotations. Therefore, we focus exclusively on MCQA and Yes/No QA annotations to create an informative anchor caption.
  • Perception Test spans various skill and reasoning domains to thoroughly evaluate VLLMs’ perception and understanding abilities. Our inspection of these evaluation dimensions reveals alignment between the semantics, physics, and memory skill areas, as well as descriptive and explanatory reasoning dimensions, with the temporal aspects of action, order, and event order. Accordingly, we limit our video selection in Perception Test to these specific pillars. Additionally, we review the question templates adopted in these areas and select video instances with question-answer pairs that support VidHal’s evaluation objectives. The specific skills and associated questions chosen are detailed in Figure 14.
  • MVBench includes twenty video understanding tasks with question-answer pairs designed to challenge the reasoning capabilities of VLLMs. Similar to the Perception Test, we identify the tasks relevant to the temporal aspects in VidHal and focus on collecting videos belonging from these tasks. The specific tasks for each aspect are presented in Figure15. We observe that MVBench contains repeated use of certain scenarios across tasks, indicated by similar question templates. To enhance caption diversity and minimize redundancy, we limit the number of examples for each unique scenario. The collected instances cover all five temporal aspects of VidHal.
  • AutoEval-Video evaluates open-ended response generation in VLLMs through questions with detailed answers across nine skill dimensions. We focus on instances related to the state transition area, specifically assessing changes in object and entity attributes. For each instance, we retain the only answers to associated questions as they act as informative, long-form captions for the video.
Incorrect Anchor Captions
A minority of videos contain anchor captions misaligned with their content, often due to noisy metadata. Such discrepancies subsequently lead to undesirable hallucinatory captions. To remove such instances, we use BLIP2 [27] to calculate frame-text matching scores across all video frames, selecting the maximum score as the representative video-text alignment score. Examples with incorrect anchor captions typically achieve low alignment scores, which are discarded as noisy instances.
LLM-based Caption Generation
We utilize GPT-4o’s [41] text processing and generation capabilities to generate an anchor caption for each selected video, based on metadata from its original public dataset source. This metadata includes QA-based annotations for TempCompass, Perception Test, and MVBench, along with long-form answers for AutoEval-Video. The anchor caption is subsequently used as input for GPT-4o to generate corresponding hallucinatory captions.
To ensure the generated hallucinatory captions meet high-quality standards, we employ a detailed prompt adopting the following strategies to guide GPT-4o’s output:
  • Aspect-specific definitions which outline the characteristics of each aspect to be varied, prompting GPT-4o to modify anchor captions accordingly.
  • Caption construction guidelines that define the structure, format, and hallucination levels required for the generated captions.
  • In-context examples to illustrate the desired form of each hallucinatory caption for each aspect.
The prompts for generating anchor and hallucinatory captions are shown in Figures 16 and 17, respectively. Definitions for each aspect are provided in Figure 18, and aspect-specific in-context examples are detailed in Figures 19 to 27. Separate in-context examples are provided for each Attribute subaspect of Shape, Size, Color, Count, and State Change to account for their distinct natures.
Figure 16: Prompts used for generating the anchor caption from (Top) QA-based annotations and (Bottom) long-form captions.
Figure 17: Prompt for generating aspect-specific hallucinatory captions based on anchor captions and in-context examples.
Figure 18: Definitions incorporated into the prompt for generating hallucinatory captions for each aspect, with separate definitions provided for each sub-aspect in the Attribute aspect.
Figure 19: In-context examples for Size sub-aspect.
Figure 20: In-context examples for Shape sub-aspect.
Figure 21: In-context examples for the Color sub-aspect.
Figure 22: In-context examples for the Count sub-aspect.
Figure 23: In-context examples for the State sub-aspect.
Figure 24: In-context examples for the Object aspect.
Figure 25: In-context examples for the Event Order aspect.
Figure 26: In-context examples for the Action aspect.
Figure 27: In-context examples for the Direction aspect.
Figure 28: Prompt template for evaluating the quality of generated captions for the GPT-4o, Gemini-1.5 Flash, and LLaMA3 (70B) models.
Figure 29: Question prompts for evaluating caption quality based on the three assessment criteria. Prompts with the placeholder <option> are applied individually to the anchor and hallucinatory captions. For question associated with order quality, <option_A> and <option_B> are replaced with the corresponding hallucinatory caption options shown to the LLMs.
Caption Quality Scoring
To identify video instances with the high quality generated captions, we utilize powerful LLMs to evaluate the quality of generated captions. The captions are assessed is based on three specific criteria:
  • Realism determines whether generated scenarios are plausible.
  • Ordering Quality evaluates whether the hallucination level ordering is appropriate.
  • Relevance ensures that deviations from the anchor caption align with the designated aspect.
Binary questions are used to evaluate captions for each criterion, assigning a score of 1 for positive responses, i.e., ”yes”, and 0 otherwise. The scores for each criterion are averaged across all models and prompts, and then summed across all criteria to produce a final quality assessment score for the generated captions of a video instance. We evaluate each set of captions using three LLMs: GPT-4o, Gemini-1.5 Flash [44], and LLaMA3 (70B) [12] along with three variants for each binary question. This ensemble of both models and prompts enhances the robustness of our evaluation.. Figures 28 and 29 provide details of the criterion-specific quality assessment queries and the prompt templates employed for each LLM. We select the top 1,000 examples with the highest quality assessment scores to construct VidHal.
Figure 30: Qualitative examples of video instances and their corresponding generated captions in the VidHal Benchmark, across the five temporal aspects.
A.3 Additional Dataset Examples
We provide additional qualitative examples of video instances and their corresponding captions in Figure 30 for each of the five temporal aspects.
B Human Validation Details
B.1 Human Validation Process
As varying hallucination levels are a distinctive feature of our benchmark, we prioritize validating the robustness of caption ordering produced by our annotation pipeline. Each anchor caption is derived from the original video metadata, making it the most accurate reflection of the video content. Our primary objective is to ensure that the ordering of hallucinatory captions aligns with human judgment. To achieve this, human annotators are shown the video instance along with both hallucinatory captions and are tasked with selecting the caption that better aligns with the video content, as illustrated in Figure 31. Each video instance is reviewed by multiple annotators, with the final human-aligned order determined through a majority vote and compared with our automatically generated order.
Figure 31: Pipeline for validating the quality of generated caption orders in VidHal. For each instance, human annotators are provided with the video and its associated hallucinatory captions. The annotators then select the caption that best aligns with the video content. The selected response is subsequently checked for consistency with the caption with lower hallucination according to our annotation process.
B.2 Misaligned Instances
Table 4 lists video instances that fail to meet the majority agreement threshold established by our annotation process. We additionally provide the corresponding human agreement scores for each instance.
Video ID Agreement Score
action_55 0.429
action_88 0
action_90 0.308
action_118 0.200
action_153 0.250
order_60 0.500
order_109 0.154
attribute_90 0.400
attribute_180 0.071
attribute_192 0.188
object_25 0.375
object_170 0
direction_188 0.400
Table 4: Video examples with generated caption orders misaligned with the human-preferred order during the quality verification process. The agreement score indicates the proportion of human respondents who select the ordering generated in our annotations.
C Evaluation Pipeline Details
C.1 Evaluation Task Prompts
Figures 32 and 33 present the prompts used for the MCQA and naive caption ordering tasks, respectively. The same prompt used for both the MCQA task and the paired questions in the relative caption ordering task.
Figure 32: Prompt template for the MCQA and relative caption ordering evaluation tasks.
Figure 33: Prompt template for the naive caption ordering evaluation task.
C.2 Relative Order Parsing
Prompting the VLLM to predict the order of captions based on their hallucinatory level in the relative caption ordering task involves asking a series of paired questions derived from different caption combinations. However, providing the model with all possible pairs at once may result in cyclic and non-transitive orderings. To address this, we present each caption pair to the VLLM in a systematically selected sequence, beginning with two paired questions. The final paired question is presented to the model to resolve inconsistencies if the multiple possible orderings can be derived from the responses to the first two paired questions. The responses across all paired questions presented to the VLLM is then parsed according to the workflow illustrated in Figure 34.
Figure 34: Decision tree for determining the final caption order based on VLLM responses to paired questions in the relative caption ordering evaluation task.

### v2

@@ 5.1 Experimental Settings
Models. We evaluated thirteen VLLMs from eight different model families, including six open-source models: VideoChat [27], LLaMA-VID [30], VideoLLaMA2 [8], VideoChat2 [28], mPLUG-Owl3 [59] and LLaVA-NeXT-Video [66], and two proprietary models: GPT-4o [42] and Gemini-1.5 [45]. These models represent a wide variety of architectural designs and training paradigms. Additionally, we included a random baseline that selects and ranks candidate options randomly.
Implementation Details. All experiments were conducted using four NVIDIA A100 40GB GPUs. The input captions in $\mathcal{Y}^{i}$ were presented in a randomized order using a fixed, predefined randomization seed across experiments. We adhered to the inference and model hyperparameters outlined in the respective original models, and employed greedy decoding during generation for a fair comparison.
@@ 5.2 Overall Results
We present the overall results of representative VLLMs in Table 2 across both MCQA and caption ordering tasks. We make three key observations from this table: I) Proprietary models demonstrate superior results compared to open-sourced models. In particular, GPT-4o achieves the best performance on all tasks, surpassing other models by significant margins. II) Larger VLLMs generally outperform smaller ones in both tasks. This result is supported by the comparison of different LLM bases for the VideoLLaMA2 and LLaVA-NeXT-Video models. III) The caption ordering task poses greater difficulty for current VLLMs than MCQA, evidenced by the larger performance margins between the VLLM models and the random baseline. Notably, VideoChat and VideoChat2 (Mistral) show slight to no improvement over the random baseline across both caption ordering tasks. This indicates that current VLLMs greatly suffer from poor fine-grained video understanding and are inclined to generate hallucinations.
Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task. The dotted lines represent the average NDCG scores across all models.
@@
We further analyzed the distribution of results for the relative caption ranking task across sub-aspects of the Attribute and Object aspects in Figure 6. While VLLMs generally maintain consistent performance across Attribute sub-aspects, they demonstrated slightly less effectiveness in inference involving Count and Color. For the Object aspect, several models performed significantly worse in Interaction Classification than in Object Recognition. This highlights the need for greater emphasis on modeling object-based interactions to bridge the gap between object recognition and interaction understanding in VLLMs.
Figure 6: NDCG scores for the Attribute (Left) and Object (Right) sub-aspects in the relative caption ranking task.
@@ 5.3 Ablation Studies
VideoChat2 (Vicuna) mPLUG-Owl3 LLaVA-NeXT -Video (7B) VideoLLaMA2 (7B)
$P_{\text{TI}}\ (\%)$ 7.3 3.0 3.7 4.0
Table 3: Percentage of VidHal instances where final predicted orders contradict individual paired-caption orders.
Transitive Robustness
Despite improving VLLM stability in caption ordering, the relative caption ranking task may reveal transitive inconsistencies between the final caption order and individual paired orders. For instance, the predicted order $\mathcal{\hat{Y}}^{i}=[\hat{y}^{i,1},\hat{y}^{i,2},\hat{y}^{i,3}]$ may be inferred when the VLLM predicts $\hat{y}^{i,1}\prec\hat{y}^{i,2}$ and $\hat{y}^{i,2}\prec\hat{y}^{i,3}$, but it may predict $\hat{y}^{i,3}\prec\hat{y}^{i,1}$ when prompted to order these two captions, contradicting the initial order. We quantify such errors by measuring the proportion of instances where $\mathcal{\hat{Y}}^{i}$ does not entail $\hat{y}^{i,3}\prec\hat{y}^{i,1}$. This is formally defined as:
$P_{\text{TI}}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\{\hat{y}^{i,1},\hat{y}^{i,3}\})=\hat{y}^{i,3}\right].$ (7)
The results, presented in Table 3, demonstrate strong consistency between individual paired caption orders and final caption orders in VLLM predictions.
Figure 7: Overlapping ratios of model predictions under single-frame and full-video inputs for correct, incorrect and overall predictions. Complete Reliance indicates that the VLLM always produces the same response for both video and single-frame inputs.
Image Prior Reliance
Prior research reveals that VLLMs often leverage image priors for reasoning [22, 3], neglecting important spatiotemporal features. This phenomenon is exemplified by a few frames having a dominant influence on response generation. To investigate this bias influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [47] to extract the most salient frame $v^{i}$ from $V^{i}$. Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input. The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings. As illustrated in Figure 7, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions. This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
@@
5.4 Further Analysis
Response Stability
The unique structure of the naive caption ordering task, distinct from existing visual understanding tasks, may lead to inadequate model responses due to potential gaps in understanding. To identify these deficiencies, we use the two metrics outlined below:
@@ 5.4 Further Analysis
Figures 8 presents the IRR and RR scores for all evaluated models. Despite being able to formulate responses with the correct structure, many VLLMs tend to produce the same caption order regardless of $V^{i}$, as reflected by their high RR scores. This is particularly evident in models with limited capacity, such as those in the 7B parameter range.
Hallucination Differentiation Sensitivity
We investigate the tendency of VLLMs to favor captions with higher hallucination over those with lower degree in the relative caption ranking task. For two captions with different hallucination levels $j,k$ where $j>k$, we introduce the following metric to quantify such hallucination misalignment cases:
@@
which reflects the proportion of cases in which the VLLM selects the caption with a higher level of hallucination $j$ over $k$. Specifically, we examine three key cases: when the most hallucinatory caption is chosen over both the lower-hallucination and anchor captions, and when the lower-hallucination caption is selected over the anchor caption. These cases are represented by $HM_{3\rightarrow 1}$, $HM_{3\rightarrow 2}$, and $HM_{2\rightarrow 1}$, respectively, with results presented in Figure 9.
Our findings show that advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9. However, two key observations emerge from our experiments: First, most VLLMs struggle to differentiate the lower hallucinatory caption from the anchor, as evidenced by the gap between $HM_{3\rightarrow 1}$ and $HM_{2\rightarrow 1}$. Second, all models exhibit high $HM_{3\rightarrow 2}$ scores, indicating difficulty in distinguishing between two hallucinatory captions with varying degrees. These results suggests that gaps in nuanced video reasoning may contribute to hallucinatory behavior in VLLMs, a challenge not addressed by existing ⟨positive, hallucinatory⟩-based evaluation methods. [29, 51, 52, 16].
@@ 6 Conclusion
In this work, we introduce the VidHal benchmark to address gaps in the video-based hallucination evaluation of VLLMs. VidHal features video instances spanning five temporal aspects. Additionally, we propose a novel caption ordering evaluation task to probe the fine-grained video understanding capabilities of VLLMs. We conduct extensive experiments on VidHal through the evaluation of thirteen VLLMs, exposing their limitations in unexpected hallucination generation. Our empirical results shed light on several promising directions for future work: e.g., incorporating a broader range of temporal features during pretraining and mitigating single-frame priors to enhance temporal reasoning. These advancements will help to address the hallucination problem in video-based VLLMs, enhancing their robustness for real-world video understanding applications.